evaluate: a^2-b^2 over a-b when a=3 and b= -4


Answer by babynarah2007(12)

Solution by EXPLAIN simplification of an expression
Your Result:
Simplify a^2-b^2overa-b when a=3 and b=-4
use simply substitution/evaluating
a^2-b^2 over a-b mean: divide a^2-b^2/a-b
so,a=3and b=-4 mean simply them to the equations
(3)^2-(-4)^2/3-(-4)
= 9-16/3+4
=-7/7
=-1 (now it's done!)
the answer is (-1)
Happy X'mas and Happy New year